OpenStudy (anonymous):

Please help!!! (2b)^-5=????

OpenStudy (anonymous):

you mean \[\left( 2b \right)^\left( -5 \right) = ?\]

OpenStudy (anonymous):

I am thinking (2)^-5 x (b)^-5= 42^-5.b^-5..not sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it can be written as \[\frac{ 1 }{ \left( 2b \right)^\left( 5 \right) }\]

OpenStudy (anonymous):

and now i say its easy to solve :)